OpenStudy (anonymous):

a square graden has a side of 22 meters how many meters of fence are needed to enclose the garden?

OpenStudy (superdavesuper):

Its a square n its side=22 meters....the fence will form the perimeter. wats the perimeter of a square equal to?

OpenStudy (anonymous):

emily_96 help!!!!

OpenStudy (anonymous):

o wait nvm

OpenStudy (anonymous):

thnx superdave super

OpenStudy (superdavesuper):

welcome....wat u have for ans?

OpenStudy (anonymous):

what u mean?

OpenStudy (superdavesuper):

what is ur answer? just want to make sure u got it right :)

OpenStudy (anonymous):

wait i havnt got it yet still doing it give me one more sec

OpenStudy (anonymous):

22x 4?

OpenStudy (anonymous):

88?

OpenStudy (superdavesuper):

that is the correct answer, good job! :)

OpenStudy (anonymous):

can u help me?

OpenStudy (superdavesuper):

got another Q? sure :)

OpenStudy (anonymous):

a chess board has an area of 100 square inches what is its perimeter??

OpenStudy (superdavesuper):

this a lil more complicated. a chess board is a square. u know its 100 sq inches, so what is the length of one side of the board? remember its a square...

OpenStudy (anonymous):

100x 4

OpenStudy (anonymous):

400?

OpenStudy (superdavesuper):

nope...wats the length of 1 side of the square board? length; not area.

OpenStudy (anonymous):

idk? :(

OpenStudy (superdavesuper):

ok assume the side of square is a, whats the area of the square?

OpenStudy (anonymous):

400 i think but no actually cause its a square its 400 so ya

OpenStudy (superdavesuper):

Nope its not 400. Will u answer this question first plz? assume the lengh of one side of the square is a, whats the area of the square?

OpenStudy (anonymous):

i really dont know thiz its a 6 th grade queetion like c mon

OpenStudy (anonymous):

i dont know

OpenStudy (anonymous):

i tried multiplying by width by length

OpenStudy (superdavesuper):

OK area of a square with side =a is.... like u just said: width by length which is the same because it is a square! so area = a*a=a^2 Got it?

OpenStudy (anonymous):

oh omg thnx

OpenStudy (superdavesuper):

So what is the length of one side of the chess board if its area is 100?

OpenStudy (anonymous):

100x 100 =1000

OpenStudy (superdavesuper):

sorry read it again...I think u got it backward...u know area n want to find length...

OpenStudy (anonymous):

dont get it bye

OpenStudy (superdavesuper):

Length of one side is 10 because 10*10=100